\begin{thebibliography}{100}

	\bibitem{Butler06} Butler, R. P., and 10 colleagues (2006) Catalog of Nearby 
Exoplanets.  {\em The Astrophysical Journal} 646: 505-522.

	\bibitem{Wright11} Wright, J. T., and 10 colleagues (2011) The Exoplanet Orbit 
Database.  {\em Publications of the Astronomical Society of the Pacific} 
123: 412-422.

	\bibitem{MayorQueloz95} Mayor, M. and Queloz, D. (1995) A Jupiter-mass companion to a 
solar-type star.  {\em Nature} 378: 355-359.

	\bibitem{Lissauer93} Lissauer, J. J. (1993) Planet formation.  {\em Annual Review of 
Astronomy and Astrophysics} 31: 129-174.

	\bibitem{KleyNelson12} Kley, W. and Nelson, R. P. (2012) Planet-Disk Interaction and 
Orbital Evolution.  {\em Annual Review of Astronomy and Astrophysics} 50: 
211-249.

	\bibitem{Capobianco11} Capobianco, C. C., Duncan, M., and Levison, H. F. (2011) 
Planetesimal-driven planet migration in the presence of a gas disk.  {\em 
Icarus} 211: 819-831.

	\bibitem{RasioFord96} Rasio, F. A. and Ford, E. B. (1996) Dynamical instabilities 
and the formation of extrasolar planetary systems.  {\em Science} 274: 
954-956. 

%

	\bibitem{WeidenschillingMarzari96} Weidenschilling, S. J. and Marzari, F. (1996) Gravitational 
scattering as a possible origin for giant planets at small stellar 
distances.  {\em Nature} 384: 619-621.

	\bibitem{FabryckyTremaine07} Fabrycky, D. and Tremaine, S. (2007) Shrinking Binary and 
Planetary Orbits by Kozai Cycles with Tidal Friction.  {\em The 
Astrophysical Journal} 669: 1298-1315.

	\bibitem{KleyNelson08} Kley, W. and Nelson, R. P. (2008) Planet formation in binary 
stars: the case of $\gamma$ Cephei.  {\em Astronomy and Astrophysics} 486: 
617-628.

	\bibitem{FordRasio06} Ford, E. B. and Rasio, F. A. (2006) On the Relation between 
Hot Jupiters and the Roche Limit.  {\em The Astrophysical Journal} 638: 
L45-L48.

	\bibitem{LinIda97} Lin, D. N. C. and Ida, S. (1997) On the Origin of Massive 
Eccentric Planets.  {\em The Astrophysical Journal} 477: 781-.

	\bibitem{AdamsLaughlin03} Adams, F. C. and Laughlin, G. (2003) Migration and dynamical 
relaxation in crowded systems of giant planets.  {\em Icarus} 163: 290-306.

	\bibitem{Chatterjee08} Chatterjee, S., Ford, E. B., Matsumura, S., and Rasio, F. A. 
(2008) Dynamical Outcomes of Planet-Planet Scattering.  {\em The 
Astrophysical Journal} 686: 580-602.

	\bibitem{JuricTremaine08} Juric, M. and Tremaine, S. (2008) Dynamical Origin of 
Extrasolar Planet Eccentricity Distribution.  {\em The Astrophysical 
Journal} 686: 603-620.

	\bibitem{Holman97} Holman, M., Touma, J., and Tremaine, S. (1997) Chaotic 
variations in the eccentricity of the planet orbiting 16 Cygni B.  {\em 
Nature} 386: 254-256.

	\bibitem{LaughlinAdams98} Laughlin, G. and Adams, F. C. (1998) The Modification of 
Planetary Orbits in Dense Open Clusters.  {\em The Astrophysical Journal} 
508: L171-L174. 

%

	\bibitem{Malmberg11} Malmberg, D., Davies, M. B., and Heggie, D. C. (2011) The 
effects of fly-bys on planetary systems.  {\em Monthly Notices of the Royal 
Astronomical Society} 411: 859-877.

	\bibitem{Boley12} Boley, A. C., Payne, M. J., and Ford, E. B. (2012) Interactions 
between Moderate- and Long-period Giant Planets: Scattering Experiments for 
Systems in Isolation and with Stellar Flybys.  {\em The Astrophysical 
Journal} 754: 57-.

	\bibitem{Kaib13} Kaib, N. A., Raymond, S. N., and Duncan, M. (2013) Planetary 
system disruption by Galactic perturbations to wide binary stars.  {\em 
Nature} 493: 381-384.

	\bibitem{Katz97} Katz, J. I. (1997) Single Close Encounters Do Not Make 
Eccentric Planetary Orbits.  {\em The Astrophysical Journal} 484: 862-.

	\bibitem{MazehShahm79} Mazeh, T. and Shaham, J. (1979) The orbital evolution of close 
triple systems - The binary eccentricity.  {\em Astronomy and Astrophysics} 
77: 145-151.

	\bibitem{Ford00b} Ford, E. B., Kozinsky, B., and Rasio, F. A. (2000) Secular 
Evolution of Hierarchical Triple Star Systems.  {\em The Astrophysical 
Journal} 535: 385-401.

	\bibitem{Naoz11} Naoz, S., Farr, W. M., Lithwick, Y., Rasio, F. A., and 
Teyssandier, J. (2011) Hot Jupiters from secular planet-planet 
interactions.  {\em Nature} 473: 187-189.

	\bibitem{ZakamskaTremaine04} Zakamska, N. L. and Tremaine, S. (2004) Excitation and 
Propagation of Eccentricity Disturbances in Planetary Systems.  {\em The 
Astronomical Journal} 128: 869-877.

	\bibitem{WuLithwick11} Wu, Y. and Lithwick, Y. (2011) Secular Chaos and the 
Production of Hot Jupiters.  {\em The Astrophysical Journal} 735: 109-.

	\bibitem{Levison98} Levison, H. F., Lissauer, J. J., and Duncan, M. J. (1998) 
Modeling the Diversity of Outer Planetary Systems.  {\em The Astronomical 
Journal} 116: 1998-2014.

	\bibitem{Nagasawa08} Nagasawa, M., Ida, S., and Bessho, T. (2008) Formation of Hot 
Planets by a Combination of Planet Scattering, Tidal Circularization, and 
the Kozai Mechanism.  {\em The Astrophysical Journal} 678: 498-508.

	\bibitem{Matsumura10} Matsumura, S., Thommes, E. W., Chatterjee, S., and Rasio, F. 
A. (2010) Unstable Planetary Systems Emerging Out of Gas Disks.  {\em The 
Astrophysical Journal} 714: 194-206. 

%

	\bibitem{MoeckelArmitage12} Moeckel, N. and Armitage, P. J. (2012) Hydrodynamic outcomes 
of planet scattering in transitional discs.  {\em Monthly Notices of the 
Royal Astronomical Society} 419: 366-376.

	\bibitem{Albrecht12} Albrecht, S., and 12 colleagues (2012) Obliquities of Hot 
Jupiter Host Stars: Evidence for Tidal Interactions and Primordial 
Misalignments.  {\em The Astrophysical Journal} 757: 18-.

	\bibitem{Sanchis-Ojeda13} Sanchis-Ojeda, R., and 21 colleagues (2013) Kepler-63b: A 
Giant Planet in a Polar Orbit around a Young Sun-like Star.  {\em The 
Astrophysical Journal} 775: 54-.

	\bibitem{FordRasio08} Ford, E. B. and Rasio, F. A. (2008) Origins of Eccentric 
Extrasolar Planets: Testing the Planet-Planet Scattering Model.  {\em The 
Astrophysical Journal} 686: 621-636.

	\bibitem{Naoz12} Naoz, S., Farr, W. M., and Rasio, F. A. (2012) On the 
Formation of Hot Jupiters in Stellar Binaries.  {\em The Astrophysical 
Journal} 754: L36-.

	\bibitem{MortonJohnson11b} Morton, T. D. and Johnson, J. A. (2011) Discerning Exoplanet 
Migration Models Using Spin-Orbit Measurements.  {\em The Astrophysical 
Journal} 729: 138-.

	\bibitem{Batygin12} Batygin, K. (2012) A primordial origin for misalignments 
between stellar spin axes and planetary orbits.  {\em Nature} 491: 418-420.

	\bibitem{Lai12} Lai, D. (2012) Tidal dissipation in planet-hosting stars: 
damping of spin-orbit misalignment and survival of hot Jupiters.  {\em 
Monthly Notices of the Royal Astronomical Society} 423: 486-492.

	\bibitem{Winn10} Winn, J. N., Fabrycky, D., Albrecht, S., and Johnson, J. A. 
(2010) Hot Stars with Hot Jupiters Have High Obliquities.  {\em The 
Astrophysical Journal} 718: L145-L149.

	\bibitem{MoorheadAdams05} Moorhead, A. V. and Adams, F. C. (2005) Giant planet migration 
through the action of disk torques and planet planet scattering.  {\em 
Icarus} 178: 517-539.

	\bibitem{Steffen12a} Steffen, J. H., and 11 colleagues (2012) Kepler constraints on 
planets near hot Jupiters.  {\em Proceedings of the National Academy of 
Science} 109: 7982-7987.

	\bibitem{Narayan05} Narayan, R., Cumming, A., and Lin, D. N. C. (2005) Radial 
Velocity Detectability of Low-Mass Extrasolar Planets in Close Orbits.  
{\em The Astrophysical Journal} 620: 1002-1009.

	\bibitem{CresswellNelson06} Cresswell, P. and Nelson, R. P. (2006) On the evolution of 
multiple protoplanets embedded in a protostellar disc.  {\em Astronomy and 
Astrophysics} 450: 833-853.

	\bibitem{Knutson13} Knutson, H.~A., and 12 colleagues (2013)
Friends of Hot Jupiters I: A Radial Velocity Search for Massive, Long-Period Companions in Hot Jupiter Systems.  {\em ArXiv e-prints} arXiv:1312.2954-.

	\bibitem{Mandell07} Mandell, A. M., Raymond, S. N., and Sigurdsson, S. (2007) 
Formation of Earth-like Planets During and After Giant Planet Migration.  
{\em The Astrophysical Journal} 660: 823-844.

	\bibitem{AlexanderPascucci12} Alexander, R. D. and Pascucci, I. (2012) Deserts and pile-ups 
in the distribution of exoplanets due to photoevaporative disc clearing.  
{\em Monthly Notices of the Royal Astronomical Society} 422: L82-.

	\bibitem{Tsiganis05} Tsiganis, K., Gomes, R., Morbidelli, A., and Levison, H. F. 
(2005) Origin of the orbital architecture of the giant planets of the Solar 
System.  {\em Nature} 435: 459-461.

	\bibitem{Ford01} Ford, E. B., Havlickova, M., and Rasio, F. A. (2001) Dynamical 
Instabilities in Extrasolar Planetary Systems Containing Two Giant Planets.  
{\em Icarus} 150: 303-313.

	\bibitem{MarzariWeidenschilling02} Marzari, F. and Weidenschilling, S. J. (2002) Eccentric 
Extrasolar Planets: The Jumping Jupiter Model.  {\em Icarus} 156: 570-579.

	\bibitem{Raymond10} Raymond, S. N., Armitage, P. J., and Gorelick, N. (2010) 
Planet-Planet Scattering in Planetesimal Disks. II. Predictions for Outer 
Extrasolar Planetary Systems.  {\em The Astrophysical Journal} 711: 
772-795. 

%

	\bibitem{ShenTurner08} Shen, Y. and Turner, E. L. (2008) On the Eccentricity 
Distribution of Exoplanets from Radial Velocity Surveys.  {\em The 
Astrophysical Journal} 685: 553-559.

	\bibitem{Zakamska11} Zakamska, N. L., Pan, M., and Ford, E. B. (2011) 
Observational biases in determining extrasolar planet eccentricities in 
single-planet systems.  {\em Monthly Notices of the Royal Astronomical 
Society} 410: 1895-1910.

	\bibitem{Sumi11} Sumi, T., and 38 colleagues (2011) Unbound or distant 
planetary mass population detected by gravitational microlensing.  {\em 
Nature} 473: 349-352.

	\bibitem{Veras09} Veras, D., Crepp, J. R., and Ford, E. B. (2009) Formation, 
Survival, and Detectability of Planets Beyond 100 AU.  {\em The 
Astrophysical Journal} 696: 1600-1611.

	\bibitem{Marois08} Marois, C., and 7 colleagues (2008) Direct Imaging of Multiple 
Planets Orbiting the Star HR 8799.  {\em Science} 322: 1348-.

	\bibitem{Dodson-Robinson09} Dodson-Robinson, S. E., Veras, D., Ford, E. B., and Beichman, 
C. A. (2009) The Formation Mechanism of Gas Giants on Wide Orbits.  {\em 
The Astrophysical Journal} 707: 79-88.

	\bibitem{Kratter10} Kratter, K. M., Murray-Clay, R. A., and Youdin, A. N. (2010) 
The Runts of the Litter: Why Planets Formed Through Gravitational 
Instability Can Only Be Failed Binary Stars.  {\em The Astrophysical 
Journal} 710: 1375-1386. 

%

	\bibitem{ScharfMenou09} Scharf, C. and Menou, K. (2009) Long-Period Exoplanets From 
Dynamical Relaxation.  {\em The Astrophysical Journal} 693: L113-L117.

	\bibitem{Ford00a} Ford, E. B., Joshi, K. J., Rasio, F. A., and Zbarsky, B. 
(2000) Theoretical Implications of the PSR B1620-26 Triple System and Its 
Planet.  {\em The Astrophysical Journal} 528: 336-350.

	\bibitem{Howard10} Howard, A. W., and 9 colleagues (2010) The Occurrence and Mass 
Distribution of Close-in Super-Earths, Neptunes, and Jupiters.  {\em 
Science} 330: 653-.

	\bibitem{Batalha13} Batalha, N. M., and 75 colleagues (2013) Planetary Candidates 
Observed by Kepler. III. Analysis of the First 16 Months of Data.  {\em The 
Astrophysical Journal Supplement Series} 204: 24-.

	\bibitem{MortonJohnson11a} Morton, T. D. and Johnson, J. A. (2011) On the Low False 
Positive Probabilities of Kepler Planet Candidates.  {\em The Astrophysical 
Journal} 738: 170-.

	\bibitem{Fressin13} Fressin, F., and 8 colleagues (2013) The False Positive Rate 
of Kepler and the Occurrence of Planets.  {\em The Astrophysical Journal} 
766: 81-.

	\bibitem{Colon12} Col\'{o}n, K. D., Ford, E. B., and Morehead, R. C. (2012) 
Constraining the false positive rate for Kepler planet candidates with 
multicolour photometry from the GTC.  {\em Monthly Notices of the Royal 
Astronomical Society} 426: 342-353.

	\bibitem{Santerne12} Santerne, A., and 8 colleagues (2012) SOPHIE velocimetry of 
Kepler transit candidates. VII. A false-positive rate of 35% for Kepler 
close-in giant candidates.  {\em Astronomy and Astrophysics} 545: A76-.

	\bibitem{Lissauer12} Lissauer, J. J., and 23 colleagues (2012) Almost All of 
Kepler's Multiple-planet Candidates Are Planets.  {\em The Astrophysical 
Journal} 750: 112-.

	\bibitem{Lissauer11b} Lissauer, J. J., and 24 colleagues (2011) Architecture and 
Dynamics of Kepler's Candidate Multiple Transiting Planet Systems.  {\em 
The Astrophysical Journal Supplement Series} 197: 8-.

	\bibitem{Fabrycky14} Fabrycky, D. C., and 18 colleagues (2012) Architecture of 
Kepler's Multi-transiting Systems: II. New investigations with twice as 
many candidates.  {\em ArXiv e-prints} arXiv:1202.6328-.

	\bibitem{ChiangLaughlin13} Chiang, E. and Laughlin, G. (2013) The minimum-mass extrasolar 
nebula: in situ formation of close-in super-Earths.  {\em Monthly Notices 
of the Royal Astronomical Society} 431: 3444-3455.

	\bibitem{PaardekooperPapaloizou09} Paardekooper, S.-J. and Papaloizou, J. C. B. (2009) On 
corotation torques, horseshoe drag and the possibility of sustained stalled 
or outward protoplanetary migration.  {\em Monthly Notices of the Royal 
Astronomical Society} 394: 2283-2296.

	\bibitem{ChatterjeeTan13} Chatterjee, S. and Tan, J. C. (2013) Inside-Out Planet 
Formation.  {\em ArXiv e-prints} arXiv:1306.0576-.

	\bibitem{BoleyFord13} Boley, A. C. and Ford, E. B. (2013) The Formation of Systems 
with Tightly-packed Inner Planets (STIPs) via Aerodynamic Drift.  {\em 
ArXiv e-prints} arXiv:1306.0566-.

	\bibitem{Kobayashi12} Kobayashi, H., Ormel, C. W., and Ida, S. (2012) Rapid 
Formation of Saturn after Jupiter Completion.  {\em The Astrophysical 
Journal} 756: 70-.

	\bibitem{HansenMurray12} Hansen, B. M. S. and Murray, N. (2012) Migration Then 
Assembly: Formation of Neptune-mass Planets inside 1 AU.  {\em The 
Astrophysical Journal} 751: 158-.

	\bibitem{Carter12} Carter, J. A., and 45 colleagues (2012) Kepler-36: A Pair of 
Planets with Neighboring Orbits and Dissimilar Densities.  {\em Science} 
337: 556-.

	\bibitem{Deck12} Deck, K. M., and 6 colleagues (2012) Rapid Dynamical Chaos in 
an Exoplanetary System.  {\em The Astrophysical Journal} 755: L21-.

	\bibitem{FangMargot13} Fang, J. and Margot, J.-L. (2013) Are Planetary Systems Filled 
to Capacity? A Study Based on Kepler Results.  {\em The Astrophysical 
Journal} 767: 115-.

	\bibitem{Ciardi13} Ciardi, D. R., and 7 colleagues (2013) On the Relative Sizes 
of Planets within Kepler Multiple-candidate Systems.  {\em The 
Astrophysical Journal} 763: 41-.

	\bibitem{VerasFord12} Veras, D. and Ford, E. B. (2012) Identifying non-resonant 
Kepler planetary systems.  {\em Monthly Notices of the Royal Astronomical 
Society} 420: L23-L27.

	\bibitem{LithwickWu12} Lithwick, Y. and Wu, Y. (2012) Resonant Repulsion of Kepler 
Planet Pairs.  {\em The Astrophysical Journal} 756: L11-.

	\bibitem{GoldreichSchlichting13} Goldreich, P. and Schlichting, H. E. (2013) Overstable 
Librations can account for the Paucity of Mean Motion Resonances among 
Exoplanet Pairs.  {\em ArXiv e-prints} arXiv:1308.4688-.

	\bibitem{Petrovich13} Petrovich, C., Malhotra, R., and Tremaine, S. (2013) Planets 
near Mean-motion Resonances.  {\em The Astrophysical Journal} 770: 24-.

	\bibitem{Ford12b} Ford, E. B., and 20 colleagues (2012) Transit Timing 
Observations from Kepler. V. Transit Timing Variation Candidates in the 
First Sixteen Months from Polynomial Models.  {\em The Astrophysical 
Journal} 756: 185-.

	\bibitem{Mazeh13} Mazeh, T., and 15 colleagues (2013) Transit Timing 
Observations from Kepler. VIII. Catalog of Transit Timing Measurements of 
the First Twelve Quarters.  {\em The Astrophysical Journal Supplement 
Series} 208: 16-.

	\bibitem{Holman10} Holman, M. J., and 40 colleagues (2010) Kepler-9: A System of 
Multiple Planets Transiting a Sun-Like Star, Confirmed by Timing 
Variations.  {\em Science} 330: 51-.

	\bibitem{Fabrycky12a} Fabrycky, D. C., and 33 colleagues (2012) Transit Timing 
Observations from Kepler. IV. Confirmation of Four Multiple-planet Systems 
by Simple Physical Models.  {\em The Astrophysical Journal} 750: 114-.

	\bibitem{Ford12a} Ford, E. B., and 51 colleagues (2012) Transit Timing 
Observations from Kepler. II. Confirmation of Two Multiplanet Systems via a 
Non-parametric Correlation Analysis.  {\em The Astrophysical Journal} 750: 
113-.

	\bibitem{Steffen12b} Steffen, J. H., and 47 colleagues (2012) Transit timing 
observations from Kepler - III. Confirmation of four multiple planet 
systems by a Fourier-domain study of anticorrelated transit timing 
variations.  {\em Monthly Notices of the Royal Astronomical Society} 421: 
2342-2354.

	\bibitem{Steffen12c} Steffen, J. H., and 17 colleagues (2012) Transit Timing 
Observations from Kepler. VI. Potentially Interesting Candidate Systems 
from Fourier-based Statistical Tests.  {\em The Astrophysical Journal} 756: 
186-.

	\bibitem{Steffen13} Steffen, J. H., and 17 colleagues (2013) Transit timing 
observations from Kepler - VII. Confirmation of 27 planets in 13 
multiplanet systems via transit timing variations and orbital stability.  
{\em Monthly Notices of the Royal Astronomical Society} 428: 1077-1087.

	\bibitem{Xie13} Xie, J.-W. (2013) Transit Timing Variation of Near-resonance 
Planetary Pairs: Confirmation of 12 Multiple-planet Systems.  {\em The 
Astrophysical Journal Supplement Series} 208: 22-.

	\bibitem{Lithwick12} Lithwick, Y., Xie, J., and Wu, Y. (2012) Extracting Planet 
Mass and Eccentricity from TTV Data.  {\em The Astrophysical Journal} 761: 
122-.

	\bibitem{WuLithwick13} Wu, Y. and Lithwick, Y. (2013) Density and Eccentricity of 
Kepler Planets.  {\em The Astrophysical Journal} 772: 74-.

	\bibitem{HaddenLithwick13} Hadden, S. and Lithwick, Y. (2013) Densities and 
Eccentricities of 163 Kepler Planets from Transit Time Variations.  {\em 
ArXiv e-prints} arXiv:1310.7942-.

	\bibitem{Dawson12} Dawson, R. I., and 6 colleagues (2012) The Photoeccentric 
Effect and Proto-hot Jupiters. II. KOI-1474.01, a Candidate Eccentric 
Planet Perturbed by an Unseen Companion.  {\em The Astrophysical Journal} 
761: 163-.

	\bibitem{Nesvorny12} Nesvorny, D., and 5 colleagues (2012) The Detection and 
Characterization of a Nontransiting Planet by Transit Timing Variations.  
{\em Science} 336: 1133-.

	\bibitem{Nesvorny13} Nesvorny, D., and 5 colleagues (2013) KOI-142, The King of 
Transit Variations, is a Pair of Planets near the 2:1 Resonance.  {\em The 
Astrophysical Journal} 777: 3-.

	\bibitem{Lissauer11a} Lissauer, J. J., and 38 colleagues (2011) A closely packed 
system of low-mass, low-density planets transiting Kepler-11.  {\em Nature} 
470: 53-58.


\end{thebibliography}

\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}
\newcommand{\Kepler}{{Kepler} }

\documentclass{pnastwo}

\url{http://TBD}
\copyrightyear{2013}
\issuedate{Issue Date}
\volume{Volume}
\issuenumber{Issue Number}

\begin{document}

\title{Architectures of Planetary Systems and Implications for their Formation}	

\author{Eric B. Ford\affil{1}{Center for Exoplanets \& Habitable Worlds, The Pennsylvania State University, 525 Davey Laboratory, State College, PA 16803, USA}\affil{2}{Department of Astronomy \& Astrophysics, The Pennsylvania State University, 525 Davey Laboratory, State College, PA 16803, USA}}

\contributor{Submitted to Proceedings of the National Academy of Sciences of the United States of America}

\maketitle

\begin{article}
\begin{abstract}
Doppler planet searches revealed that many giant planets orbit close to their host star or in highly eccentric orbits.  
These and subsequent observations inspired new theories of planet formation that invoke gravitation interactions in multiple planet systems to explain the excitation of orbital eccentricities and even short-period giant planets.
Recently, NASA's Kepler mission has identified over 300 systems with multiple
transiting planet candidates, including many potentially rocky
planets.  Most of these systems include multiple planets with sizes
between Earth and Neptune and closely-spaced orbits.  These systems
represent yet another new and unexpected class of planetary systems
and provide an opportunity to test the theories developed to explain
the properties of giant exoplanets.  Presently, we have limited knowledge about such planetary systems, mostly about their sizes and orbital periods.  
With the advent of long-term, nearly continuous monitoring by Kepler, the method of transit
timing variations (TTVs) has blossomed as a new technique for characterizing the gravitational
effects of mutual planetary perturbations for hundreds of planets.  TTVs can 
provide precise (but complex) constraints on planetary masses,
densities and orbits, even for planetary systems with faint host
stars.  In the coming years, astronomers will translate TTV observations
into increasingly powerful constraints on the formation and orbital evolution of
planetary systems with low-mass planets.  
Between TTVs, improved Doppler surveys, high-contrast imaging campaigns and microlensing surveys, astronomers can look forward to a much better understanding of planet formation in the coming decade.   
\end{abstract}

\keywords{exoplanets | planet formation | Kepler}

\abbreviations{AMD, angular momentum deficit; AU, astronomical unit; MMR, mean motion resonance; RV, radial velocity; STIPS, short-period tightly-packed inner planetary systems; TTVs, transit timing variations}


\dropcap{P}rior to the discovery of exoplanets, astronomers fine tuned theories of planet formation to explain detailed properties of the solar system.  The discovery of exoplanets has significantly increased our appreciation for the diversity of planetary systems in nature.  In this article we review the current understanding of the late-stages of evolution of planetary systems, focusing on observational constraints from radial velocity and transit observations.

\section{Hot Jupiters}
Radial velocity (RV) surveys have discovered over 400 planets, most with masses larger than that of Jupiter (http://www.exoplanets.org;  \cite{Butler06,Wright11}.  Many of the early RV discoveries were ``hot-Jupiters'', planets with orbital periods of up to several days and masses comparable to that of Jupiter or Saturn (e.g., \cite{MayorQueloz95}).  As the timespan of observations has increased, the median orbital period of RV-discovered planets has steadily increased to more than a year.  Now, we know that hot-Jupiers are a relatively rare outcome of planet formation.  Nevertheless, their existence and their orbital properties provide important clues to the planet formation process.  

\subsection{Disk Migration}
Prior to the discovery of hot-Jupiters, planet formation theories had been focused on explaining properties of the solar system
\cite{Lissauer93}.  The large masses of hot-Jupiters imply a substantial gaseous component and therefore rapid formation, before the protoplanetary disk is dispersed.  {\em In situ} formation of the rocky cores of hot-Jupiters is problematic due to the high temperature and low surface density of the disk so close to their host star.  Therefore, theorists explain hot-Jupiters starting with the formation of a rocky core at larger separations from the host star, followed by accretion of a gaseous envelope and migration to their current location.  The mechanism for migration is less clear.  There are two broad classes of models: a gradual migration through a disk \cite{KleyNelson12,Capobianco11} or the excitation of a large eccentricity followed by tidal circularization \cite{RasioFord96,WeidenschillingMarzari96,FabryckyTremaine07}.  In principle, planets could migrate through either a gaseous protoplanetary disk or a planetesimal disk.  If giant planet cores form beyond the ice line, then planetesimal disks would rarely be massive enough to drive the large-scale migration needed to form a hot-Jupiter.  On the other hand, a gaseous protoplanetary disk could power a rapid migration, perhaps too rapid \cite{KleyNelson08}.  It is unclear how the planets would avoid migrating all the way into the host star or why the migration would be halted to leave planets with orbital periods of $\sim$2-5 days \cite{FordRasio06}.  The apparent pile-up of hot-Jupiters with orbital periods of a few days could be the result of censoring, i.e., those planet that continued to migrate closer to their host star were either accreted onto the star, destroyed or reduced in mass due to stellar irradiation and mass loss.  Even in this case, a stopping mechanism must be invoked to produce the observed giant planets with orbital periods beyond $\sim$7 days, since tides rapidly become inefficient with increasing orbital separations.  

\subsection{Eccentricity Excitation plus Tidal Circularization}

Unlike disk migration, eccentricity excitation followed by tidal circularzation naturally explains the ``pile-up'' of hot-Jupiters at orbital periods of 2-7 days due to the rapid onset of tidal effects.  The large eccentricities required to initiate circularization could be generated in a variety of ways.  
The simplest scenario is planet-planet scattering, as it requires only one additional massive planet \cite{RasioFord96}.  In the case of two planets and no additional perturbers, the initial ratio of semi-major axes must be small enough to permit close encounters.  Such scenarios may arise naturally for giant planets due to rapid mass growth.  Alternatively, a system with more than two planets \cite{WeidenschillingMarzari96,LinIda97,AdamsLaughlin03,Chatterjee08,JuricTremaine08} will naturally approach instability on a much longer timescale.  Even for a simple three planet system, the timescale until close encounters can easily exceed ten million years \cite{Chatterjee08}, by which time the protoplanetary disk will have dissipated.  

Alternatively, a system of two (or more planets) may become unstable due to an external perturbation, such as secular interactions with a binary companion \cite{Holman97} or a stellar flyby \cite{LaughlinAdams98}.  This can lead to strong planet scattering, often following a prolonged phase of weaker interactions \cite{Malmberg11,Boley12,Kaib13}.  
Each close encounter between giant planets leads to a small perturbation to their orbits \cite{Katz97}.  Thus, it is typically a series of close encounters that excites the two planets' orbital eccentricities until one planet's pericenter is small enough to initiate tidal circularizaiton.  In practice, planetary systems that form two giant planets may well form additional massive planets, leading to a series of planet-planet scattering events and substantially increasing the probability for one to achieve a pericenter of just a few stellar radii.  

Another possible mechanism for eccentricity excitation involves secular (i.e., long-term) perturbations by one or more distant bodies (e.g., more planets, a brown dwarf or binary stellar companion).  If there is a large mutual inclination between the inner planet and the outer companion, then large eccentricities are possible, even for systems with large orbital period ratios \cite{MazehShahm79,Holman97,Ford00b,Naoz11}.  This could be particularly relevant for planets orbiting one member of a binary (or higher multiple) star system, even though the orbital period of the two stars is often much larger than the orbital period of the planet.  

For small mutual inclinations, exciting an eccentricity large enough to trigger tidal circularization requires a substantial angular momentum deficit (AMD) and a series of planets that serve to couple the inner giant planet to outer planets which are more likely to have a significant initial AMD \cite{ZakamskaTremaine04}.  
If the planets are widely spaced, then it is possible to construct initial conditions that lead to the inner planet's pericenter dropping to only a few stellar radii \cite{WuLithwick11}.  However, for more typical initial conditions, the secular interactions lead to close encounters between planets that result in collisions and/or ejections via planet-planet scattering \cite{LinIda97,Levison98,Chatterjee08,Nagasawa08,Matsumura10,MoeckelArmitage12}.  

\subsection{Distinguishing between Hot-Jupiter Formation Models}

In practice, nature may provide multiple migration mechanisms for forming hot-Jupiters.  For many years, observations provided little data to help distinguish between even the two broad classes of migration models.  The major breakthrough was the measurement of the Rossiter-McLaughlin effect for many transiting hot-Jupiters.  When the planet passes in front of a rotating star, the apparent radial velocity is perturbed due to the planet blocking a portion of the star that is rotating towards or away from the observer.  The Rossiter-McLaughlin signature measures the angle between the star's rotational angular momentum and the planet's orbital angular momentum (after projecting both onto the sky plane).  While many systems are well-aligned, a significant fraction of hot-Jupiters are severely misaligned \cite{Albrecht12}.  Misaligned systems, including nearly polar and even retrograde configurations \cite{Sanchis-Ojeda13}, arise naturally in scenarios that include eccentricity excitation plus tidal circularization \cite{Nagasawa08,Naoz11}.  While planet scattering and secular perturbations will only produce hot-Jupiters for a few percent of planetary systems \cite{FordRasio08,Naoz12}, this is consistent with the rate of hot-Jupiters observed, or at least a substantial fraction of them.  

Astronomers have begun to attempt to deconvolve the distribution of Rossiter-McLaughlin measurements into a mixture of systems formed through disk migration, planet scattering and secular perturbations \cite{MortonJohnson11b}.  However, we caution that there may be systematic biases in the outputs of such analyses, due to uncertainties in the treatment of tidal circularization.  Another potential confounding factor is the possibility of a primordial star-disk misalignment \cite{Batygin12,Lai12}.  Finally, we caution that an apparent trend of obliquity with stellar temperature \cite{Winn10} calls into question even qualitative predictions of tidal theory.  While the details remain unclear, the Rossiter-McLaughlin observations demonstrate that simple disk migration is inadequate to explain all hot-Jupiters.  Of course, these observations do not preclude disk migration from having operated in systems that were later sculpted by planet scattering or secular perturbations.  Indeed, planet scattering is more efficient at forming hot-Jupiters, if migration were to bring planets to ~1 AU prior to scattering \cite{MoorheadAdams05} than if scattering commenced at several AU.  

Another key observational result is the realization that hot-Jupiters are seldom accompanied by additional planets close to their host star.  This was foreshadowed by radial velocity planet searches \cite{Wright11} and dramatically confirmed by \Kepler observations of transiting hot-Jupiters \cite{Steffen12a}, as these provide precise constraints on both small planets with orbital periods of weeks to months (via photometry) and low-mass planets in or near mean-motion resonances (via transit timing variations).  This result is consistent with the broad predictions of hot-Jupiter formation via eccentricity excitation plus tidal circularization, but in stark contrast to the predictions of disk migration models \cite{Narayan05,CresswellNelson06}.  Thus, the isolation of hot-Jupiters suggests that there is a strong upper limit to the fraction of hot-Jupiters formed via disk migration.  
Recent radial velocity follow-up of systems with hot-Jupiters has found long-term radial velocity accelerations in roughly half of the surveyed hot-Jupiters \cite{Knutson13}.  The location of the second-closest planet in systems with hot-Jupiters bolsters the hypothesis that hot-Jupiters may frequently commence scattering while at an orbital distance $\sim$1 AU.  

\subsection{Summary of Hot-Jupiter Formation}
In summary, planet formation from a gaseous disk likely leads to forming many planets on low-eccentricity orbits.  Initially, close encounters lead to collisions and increasing planet masses.  Once the planets become massive enough to eject bodies from the gravitational potential well of the star, ejections become more common.  The recoil from scattering planets leads to eccentricity growth of giant planets, especially in the outer regions of the planetary system.  The AMD of planets that effectively eject smaller bodies is redistributed among all the remaining planets of a planetary system.  This leads to further close encounters and collision or ejections, depending on the masses and distances of the planets involved.  This process repeats, gradually thinning the planetary system, so that the remaining planets have masses and spacings that result in an instability timescale comparable to the age of the planetary system.  In a small fraction of systems, either the chaotic interactions of a multi-body system and/or the secular interactions of highly inclined system lead to the innermost giant planet passing close enough to the host star that tidal interactions circularize its orbit, leading to the formations of a hot-Jupiter.  While the future hot-Jupiter is circularizing, it cleans out the inner solar system by scattering any rocky planets in the inner planetary system into the star or the outer regions of the planetary system \cite{Mandell07}.

\section{Giant Planets near Snow Lines}

As RV surveys gained sensitivity to giant planets with greater orbital periods, a second population of giant planets was identified at  orbital periods from $\sim$300 days to $\sim$4 years \cite{Wright11}.  While the number of RV-discovered planets per decade in orbital period decreases for greater orbital periods, the true rate for this population could remain constant or even increase since RV surveys are incomplete for greater orbital periods.  Giant planets with orbital periods $\sim$year may well have formed further out in the disk and migrated to their current location.  The distribution for orbital separation of this population appears to peak near the predicted location of the snow line.  One possible interpretation is that many giant planets migrate to near the water snow line, i.e., the location in the disk where the solid surface density increases due to condensation of H$_2$O ice.  Since the snow line affects the formation of planetesimals, migration of giant planets towards the snow line could be accommodated by a variety of migration models, including migration through a gaseous disk, migration via planetesimal scattering or even via scattering of multiple planets or planetary cores.  

At orbital periods intermediate between that of the hot-Jupiters and giant planets near the snow line, RV surveys have revealed a paucity of giant planets with orbital periods from $\sim$10-200 days.  One possible interpretation is that giant planets typically migrate through this range of periods quickly.  For migration to strand these planets, the disk would need to dissipate during the brief window of time when the planet is passing through these intermediate orbital distances.  Possible mechanisms include photoevaporation of a gaseous disk \cite{AlexanderPascucci12} or the migration of a massive (and potentially distant) planet through a mean motion resonance exciting its eccentricity and destabilizing the planetesimal disk (e.g., Nice model)\cite{Tsiganis05}.  Either model requires some degree of tuning of parameters to strand a planet at intermediate orbital periods.  In principle, this could help explain the low abundance of giant planets at intermediate periods relative to either shorter or longer orbital periods.  

\section{Long-Period Giant Planets}

\subsection{Radial Velocity}
RV surveys have begun to discover giant planets with orbital distances ranging from a few to several AU \cite{Wright11}.  Many of these giant planets have significant eccentricities, suggesting that they may have previously ejected planets from their host planetary system.  Several studies have shown that planet-planet scattering naturally produces a broad distribution of eccentricities consistent with that observed \cite{Ford01,MarzariWeidenschilling02,MoorheadAdams05,Chatterjee08,FordRasio08,JuricTremaine08,Raymond10,Matsumura10}.  While this finding is robust to the choice of initial conditions, it also implies that it may not be possible to invert the observed eccentricity distribution to infer the architectures of young planetary systems shortly after the dissipation of the proto-planetary disk.  Thus, testing models of planet formation will require more detailed observations and comparisons.  
For example, the mass-period-eccentricity distribution may provide additional clues \cite{Raymond10}, if properly interpreted to account for observational selection effects.  Based on radial velocity observations, there appears to be a trend for more massive planets to have lower eccentricities than less massive planets \cite{FordRasio08}.  This is opposite the prediction of the simplest planet-planet scattering models.  However, the apparent trend could be easily accommodated if planet scattering often occurs while there is still a significant planetesimal disk which could damp eccentricities following the final scattering.  In this scenario, systems with more massive scattering planets would be less affected by the planetesimal disk and retain larger eccentricities than systems in which the planet-planet scattering involved less massive planets.  Thus, the mass-period-eccentricity correlation may provide a path towards constraining the late stage evolution of planetary systems.  However, one should keep in might that eccentricity measurements have a positive bias and the size of the bias increases for planets with radial velocity amplitudes small relative to the measurement precision \cite{ShenTurner08,Zakamska11}.  

Even more powerful constraints on planet formation are likely to come from the architecture and dynamical state of multiple planet systems (e.g., resonances, near resonances, amplitudes of secular modes).  
%
%
%
Unfortunately, characterizing the architectures of outer planetary systems is likely to prove significantly more difficult than planets with orbital periods of a few years or less.  Only several dozen stars have been observed at high precision long enough to characterize a Jupiter-clone.  Even for these targets, Saturn, Uranus and Neptune would appear only as small long-term trends given the duration of radial velocity planet searches.  Since an observed long-term acceleration might be due to a combination of multiple long-period planets, radial velocity surveys are more useful for placing limits on what planets with periods of decades or more are not present than for characterizing outer planetary systems.  

\subsection{Microlensing}
While direct imaging planet searches are just beginning to detect planets, microlensing surveys suggest that there is a substantial population of giant planets that has not been detected by radial velocity or transit searches \cite{Sumi11}.  These planets could either reside at large orbital separations from their host star or be free-floating after planet scattering ejected them from their original parent star.  In practice, both types of planets likely contribute to the microlensing population \cite{Veras09}.  Direct imaging searches may help disentangle this degeneracy.  

\subsection{Direct Imaging}
Recently, direct imaging planet searches have begun to detect low-mass companions with separations of tens to hundreds of AU (e.g., HR 8799)\cite{Marois08}.\footnote{It is often unclear whether these should be considered planets or brown dwarves, given our lack of knowledge of their mass or formation mechanism.  Here, we refer to all such companions as planets for simplicity, irrespective of their actual mass or formation history.}    
It is unclear whether these planets formed near their current separations or formed closer to their host star and migrated to their current location \cite{Dodson-Robinson09,Kratter10}.  Migration via a gas disk would require an unusually large disk mass at large separations.  This is an even more severe issue for planetesimal-driven migration.  Another possibility is that these planets may have migrated via planet-planet scattering, if these stars host at least one planet even more massive that the planet(s) observed at large separations \cite{ScharfMenou09,Veras09}.  In systems with extensive disks, the outer disk may have raised the planet's pericenter so that it decoupled from the inner regions of the planetary system.  Alternatively, for the young stars typically targeted by direct imaging searches, it is expected that we will detect planets that are still be in the process of being ejected from their star \cite{Veras09}.  

A final complication is that planets in wide orbits ($\ge$100 AU) could have formed around another star and exchanged into their present orbital configurations \cite{Malmberg11,Boley12}.  This does not offer an easy explanation for giant planets at very large separations, since typical exchange interactions are rapid and the ratio of the planet's final and initial semi-major axes only changes by a factor comparable to the ratio of host star masses.  Thus, exchanges may be important for explaining individual planetary systems (e.g., PSR 1620-26)\cite{Ford00a}, but are unlikely to explain the overall frequency of planets on very wide orbits.  

\section{Super-Earths \& Mini-Neptunes}

Over the past several years, radial velocity surveys began to discover planetary systems with Neptune and super-Earth-mass planets at short orbital periods \cite{Howard10}.  NASA's Kepler mission has demonstrated that such planets are much more common than giant planets and that most solar-type stars harbour a sub-Neptune-size planet \cite{Batalha13}.  Of more than 3,600 planet candidates identified by Kepler, only a small fraction have been dynamically confirmed (i.e., by radial velocity observations or transit timing variations).  However, detailed analyses of possible astrophysical false positives demonstrate that the average rate of false positives is sufficiently low that one can already begin to study the population of planet candidates as a whole \cite{MortonJohnson11a,Fressin13}, as long as one keeps in mind that  some subsets of planet candidates (e.g., giant planets with orbital periods less than $\sim~2$ days or $\sim~10-100$ days) have a higher rate of false positives \cite{Colon12,Santerne12}.  Fortunately, other important subsets, such as Neptune-size planets, can be shown to have a smaller false positive rate \cite{MortonJohnson11a,Fressin13}.  One very important subset of highly reliable planet candidates is those in systems with multiple transiting planet candidates \cite{Lissauer12}.
These can be considered planets with a confidence similar to that of planets discovered by radial velocity searches.  This allows one to conclude that planets with sizes between that of Earth and Neptune and orbital periods of weeks to months are the typical outcome of planet formation around solar-type stars.

\section{Short-Period Tightly-packed Inner Planetary Systems}

Beyond the sheer abundance of sub-Neptune-size planets, one particularly striking discovery is the abundance of systems with multiple transiting planets \cite{Lissauer11b,Fabrycky14}.  The typical system contains transiting planets with short orbital periods, ranging from $\sim$1-100 days.  The decrease in frequency at smaller orbital periods is real, while the decrease at large orbital periods may be at least partially due to detection biases, since both the geometric transit probability is lower and the signal to noise is smaller due to the smaller number of transits observed by Kepler.  
Such systems are typically tightly-packed, meaning that the orbital periods of planets transiting a common star are correlated with one another and not consistent with being drawn independently from a common orbital period distribution function.  While orbital periods range from hours to hundreds of days, the distribution of orbital period ratios is broadly peaked from $\sim$1.3 to 3.  Therefore, we refer to these systems as Short-Period Tightly-packed Inner Planetary Systems (STIPS).  

While theorists have developed multiple mechanisms for migrating giant planets from beyond the snow line to near their present location, the lower masses of sub-Neptune-size planets mean that {\em in situ} formation can not be dismissed so easily \cite{ChiangLaughlin13}.  The mass of most such planets is likely dominated by rock, ices or water, but not gas, even for low-density, Neptune-size planets.  Therefore, these planets need not accrete a substantial rocky core before the protoplanetay disk is cleared.  Indeed, it is conceivable that many of these atmospheres are primarily the result of outgassing, rather than accretion of gas from the disk.  The combination of lower masses and the possibility of a longer timescale for formation makes {\em in situ} formation worthy of serious consideration.  

Of course, just because theory does not prohibit in situ formation, one should not assume that planets in STIPS did not undergo significant migration.  Since sub-Neptune-size planets would not clear a gap, they are expected to migrate through a given gaseous disk more rapidly than giant planets.  While early work suggested that planetary migration would accelerate as planets spiral closer to their host stars, recent studies incorporating more realistic disk models predict material may accumulate or become trapped at orbital distances that are not dissimilar from the location of Kepler's sub-Neptune-size planets \cite{PaardekooperPapaloizou09}.  
Multiple mechanisms have been proposed for for setting the overall scale for the orbital distances of STIPS, including the transition to an MRI-active inner region \cite{ChatterjeeTan13} and the silicon sublimation front \cite{ChiangLaughlin13,BoleyFord13}.  Once one planet has formed at small distances, it can perturb the disk structure so that additional planets form or become trapped in resonances, not far behind \cite{Kobayashi12}.  Subsequent dynamical evolution can lead to spreading out of planetary systems \cite{HansenMurray12} and disruption of resonances.  

\subsection{Orbital Spacing of Planets in STIPS}

While dynamical stability prohibits systems of two giant planets with a period ratio less than $\sim$1.3 (excluding small islands of stability near mean motion resonances), Kepler's systems of small and presumably lower-mass planets can remain stable even for much tighter spacings such as $\sim$1.17 (Kepler-36)\cite{Carter12,Deck12}.  
The relative frequency of planetary systems with one, two and three transiting planets implies that there is a population of STIPS with low mutual orbital inclinations, since systems with high mutual inclinations would rarely result in multiple transiting planets \cite{Lissauer11b}.  Adding information about the ratios of transit durations shows that these systems also have low eccentricities \cite{FangMargot13,Fabrycky14}.   
The sizes of adjacent transiting planets in STIPS are observed to be correlated with each other \cite{Ciardi13}.  
These observations are qualitatively consistent with both models of {\em in situ} formation and formation at larger separations followed by migration to their currently observed orbits.  

While Kepler detects STIPS with two transiting planets more often than higher multiplicity systems, most of this trend is a detection effect due to the geometric detection probability.  Therefore, the typical outcome of planet formation appears to include forming multiple, tightly-packed, sub-Neptune-size planets with similar sizes and orbital separations.  Most such planetary systems are not in mean motion resonances (MMRs)\cite{VerasFord12}, but a significant fraction of such systems have ratios of orbital periods just slightly larger than that of a first-order MMR (typically 2:1, 3:2, or 4:3)\cite{Fabrycky14}.  This presents a significant challenge to planet formation theories.  Clearly, MMRs are playing a role in the planet formation process.  However, a slow and smooth migration would be expected to trap planets in mean motion resonances.  This has led some theorists to propose that pairs of planets are first trapped into MMRs and then dissipation causes the planets to be removed from resonance.  However, the amount of dissipation required to match the observed period ratios is so large that it implies uncomfortably large tidal quality factors \cite{LithwickWu12} or rapid eccentricity dissipation timescales \cite{GoldreichSchlichting13}.  An alternative model which invokes mass growth rather than migration or dissipation appears to require masses much larger than is plausible given the typical planet sizes \cite{Petrovich13}.  Preliminary work suggests that  interactions with a planetesimal disk may be important for removing planets from MMRs and leaving them with the orbital period ratios similar to those observed by Kepler. 

\subsection{Transit Timing Variations in STIPS}

In the case of a single star and a single planet, the orbit is strictly Keplerian and transit times follow a linear ephemeris.  In planetary systems with additional planets or stellar companions, the mutual gravitational interactions cause the transit (or eclipse) times to deviate from a strict periodicity.  The differences between the actual times of transit and the times predicted by the best-fit linear ephemeris are known as transit timing variations (TTVs).  
In many of the STIPS with planets near a MMR, Kepler has measured significant TTVs \cite{Ford12b,Mazeh13}.  
Even in cases with low TTV signal-to-noise or a TTV timescale significantly greater than the timescale of observations, TTVs can be used to confirm the planetary nature of the planet candidates \cite{Holman10,Fabrycky12a,Ford12a,Steffen12b,Steffen12c,Steffen13,Xie13} and to provide constraints on planet masses and eccentricities \cite{Lithwick12,WuLithwick13,HaddenLithwick13}.  While TTVs (presumably due to non-transiting planets) are often detected for planets transiting stars with no other known transiting planet candidates \cite{Ford12a,Mazeh13}, only in rare cases do these yield unique orbital solutions \cite{Dawson12,Nesvorny12,Nesvorny13}.  Nevertheless, simply observing that the frequency of detectable TTVs increases significantly for systems with multiple transiting planet candidates \cite{Ford12a,Xie13} provides information about the frequency of non-transiting planets in such systems and thus the frequency of closely-spaced planets and the distribution of mutual orbital inclinations.  

The realization that STIPS systems are typical raises new questions about the nature and formation of this new class of planets  with sizes between that of Earth and Neptune and the architecture of STIPS. 
The most basic questions have to do with the masses, densities and compositions of such planets. 
Since many STIPS exhibit significant TTVs, these systems provide a critical information for characterizing the masses and eccentricities of small planets \cite{Lissauer11a,Carter12}.  Over the next several years, detailed TTV analyses of several specific systems will provide precise masses for small planets.  Additionally, a systematic study of systems with TTVs is likely our best tool for characterizing the masses and densities of a large sample of small planets and thus the planet mass-radius relationship more generally.

\begin{acknowledgments}
E.B.F. acknowledges many valuable discussions with Aaron Boley, Sourav Chatterjee, Dan Fabrycky, Robert Morehead and Darin Ragozzine, as well as the entire Kepler science team and particularly the Kepler multi-body and transit timing variations working groups.  
This material is based on work supported by the National Aeronautics and Space Administration under grants NNX08AR04G and NNX12AF73G issued through the Kepler Participating Scientist Program and grant NNX09AB35G issued through the NASA Origins of Solar Systems program.  
E.B.F. acknowledges the support of the University of Florida, The Pennsylvania State University, the Penn State Center for Exoplanets and Habitable Worlds and the Institute for CyberScience.  
\end{acknowledgments}

\end{article}

\end{document}